# Crossings and diffusion in Poisson driven marked random connection models

### v1

@@ Abstract
Motivated by applications to stochastic homogenization, we study crossing statistics in random connection models built on marked Poisson point processes on $\mathbb{R}^{d}$. Under general assumptions, we show exponential tail bounds for the number of crossings of a box contained in the infinite cluster for supercritical intensity of the point process. This entails the non-degeneracy of the homogenized diffusion matrix arising in random walks, exclusion processes, diffusions and other models on the RCM. As examples, we apply our result to Poisson–Boolean models and Mott variable range hopping random resistor network, providing a fundamental ingredient used in the derivation of Mott’s law. The proof relies on adaptations of Tanemura’s growth process, the Duminil-Copin, Kozma and Tassion seedless renormalisation scheme, Chebunin and Last’s uniqueness, as well as new ideas.
@@ 1.1 Crossings and transport
Let us say that a random graph has a good crossing statistics if with high probability, for some $\kappa>0$, the subgraph given by the edges with conductances larger than $\kappa$ has at least order $\ell^{d-1}$ vertex-disjoint crossings of a box of side length $\ell$ in any given direction. Then, roughly, good crossing statistics ensure the non-degeneracy of $D$ in the 2-scale convergence approach of [Zhikov06, Faggionato23a] (at the basis of [Faggionato25, Faggionato22, FC]). For the supercritical Bernoulli percolation cluster (with unit conductances) the good crossing statistics were first proved by Kesten (cf. [Kesten82]*Chapter 11). In particular, the supercritical Bernoulli percolation cluster with unit or, more generally, positively lower bounded conductances is an effective graph to model transport in disordered conductors.
@@ 1.2 Main results
In this work we first aim at proving the above crossing property for more general random infinite clusters. We consider a random connection model (RCM) on a marked homogeneous Poisson point process (PPP) with density $\rho$ on ${\mathbb{R}}^{d}$, with i.i.d. marks, whose connection function $\varphi$, specifying the probability to have an edge between sites $x\not=y$, depends on $y-x$ and on the marks of $x$ and $y$ (see Section 2 for a formal definition). When the connection function is an indicator (as in the Poisson–Boolean model), the insertion of the edge is deterministic and fully determined by the marked PPP. We point out that the RCM on a marked PPP is also widely used in the context of scale-free graphs (see e.g. [Gracar22]).
Under suitable assumptions on $\varphi$ (cf. Assumptions 2.3), we prove the following. Assume that, for some $\lambda>0$, the RCM has a unique infinite cluster for densities $\rho>\lambda$. Then (cf. Theorem 2.4), for all $\rho>\lambda$ there exist positive constants $c,c^{\prime}$ such that for all $\ell\geq 1$ the unique infinite cluster of the RCM has at least $c\ell^{d-1}$ left-right (LR) vertex-disjoint crossings of the box $\Lambda_{\ell}:=[-\ell,\ell]^{d}$ with probability at least $1-e^{-c^{\prime}\ell^{d-1}}$. In many applications, it is known that $\lambda$ can be taken equal to the critical density of the model. Moreover, in [Chebunin24], the uniqueness of the infinite cluster is reduced to checking a simple irreducibility condition. We also point out that Theorem 2.4 extends to a much larger class the results of [Faggionato21] obtained for RCM on marked PPP whose connection function is an indicator function satisfying somewhat restrictive monotonicity properties.
@@
1.3 Proof outline
We conclude with a rough outline of the proof strategy of the crossing result, Theorem 2.4, emphasising some technical novelties and difficulties.
Firstly, by a classical argument going back to [Chayes86], it is enough to prove good crossing statistics in a quasi-two-dimensional setting. Then, LR crossings are constructed by studying a suitable growth process (see Section 6). It is related to the one introduced in [Tanemura93]*Section 4 for the blob model (i.e. the Poisson–Boolean model with a deterministic radius). However, we were led to extend and adapt Tanemura’s process, in order to recover crossings contained in the infinite cluster.
@@ 1.3 Proof outline
The above complication of ‘bad’ marks also impacts the dynamical renormalisation performed in Section 7 and illustrated in Figure 1. Indeed, the necessity to find ‘good’ marks progressively offsets renormalisation, imposing correcting the positioning at each step. Furthermore, preserving sufficient independence throughout the dynamical exploration in the continuum setting also requires some care in devising the renormalization.
@@ 2.1 Random connection model crossings
Definition 2.1 (Graph $G_{d}$).
Given $(\omega,\underline{u})\in\Omega_{*}\times[0,1]^{J}$ we denote by $G_{d}(\omega,\underline{u})$ the undirected graph on ${\mathbb{R}}^{d}$ with vertex set $\hat{\omega}=\{x_{1},x_{2},\dots\}$ and edges $\{x_{i},x_{j}\}$ where $1\leq i<j$ and $u_{i,j}\leq\varphi\big{(}(x_{i},m_{i})),(x_{j},m_{j}))$. The graph $G_{d}(\omega,\underline{u})$ defined a.s. on the space $\Omega\times[0,1]^{J}$ with probability $P_{\rho,\nu}\otimes Q$ is called Poisson driven marked random connection model (RCM) with intensity $\rho$, mark distribution $\nu$ and connection function $\varphi$. We denote this random graph by RCM$(\rho,\nu,\varphi)$.
Briefly, the RCM$(\rho,\nu,\varphi)$ is obtained as follows: sample a homogeneous PPP $\xi$ on ${\mathbb{R}}^{d}$ with intensity $\rho$, then mark points $x$ in $\xi$ by i.i.d. random variables $m_{x}$ with value in ${\mathbb{M}}$ and common distribution $\nu$ independently from the rest, then insert an edge between $x\not=y$ in $\xi$ with probability $\varphi\big{(}(x,m_{x}),(y,m_{y})\big{)}$ independently for any pair $\{x,y\}$ in $\xi$ and independently from the rest.
@@
${\mathbb{S}}_{\ell}^{-}:=\{x\in{\mathbb{S}}_{\ell}\,:\,x_{1}<-{\ell}\}\,,\qquad{\mathbb{S}}_{\ell}^{+}:=\{x\in{\mathbb{S}}_{\ell}\,:\,x_{1}>{\ell}\}\,,\qquad\Lambda_{\ell}:=[-{\ell},{\ell}]^{d}\,.$
@@
Assumption 2.3(iv) is known to hold in several models. Chebunin and Last [Chebunin24] (see there for a detailed discussion on previous results) provide a very general criterion implying Assumption 2.3(iv). To state it we introduce, for $n\in{\mathbb{N}}_{+}$, the $n$–th convolution of $\varphi$ given by the function $\varphi^{(n)}:({\mathbb{R}}^{d}\times{\mathbb{M}})^{2}\to[0,1]$ defined recursively by
$\varphi^{(1)}(t_{1},t_{2}):=\varphi(t_{1},t_{2})\qquad\text{ and }\qquad\varphi^{(n+1)}(t_{1},t_{2}):=\int\varphi^{(n)}(t_{1},t)\varphi(t,t_{2})(\lambda\mathcal{L}\otimes\nu)(dt)\,.$ (1)
@@
${\mathbb{P}}\big{(}t\text{ is connected to }t^{\prime}\text{ in }G(\xi\cup\{t,t^{\prime}\})\big{)}>0\text{ for $(\lambda\mathcal{L}\otimes\nu)^{\otimes 2}$--a.a.\ $(t,t^{\prime})\in({\mathbb{R}}^{d}\times{\mathbb{M}})^{2}$}\,,$ (3)
@@
Suppose the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. Then for any $\rho>\lambda$ there exist $c_{1},c_{2}>0$ such that, for all $\ell\geq 1$, it holds
@@
We call length of a LR crossing the number of its edges. By a large deviation estimate for Poisson random variables, for $\ell\geq 1$ the event $\sharp(\hat{\omega}\cap\Lambda_{\ell})\leq 2\rho(2\ell)^{d}$ has $P_{\rho,\nu}$–probability at least $1-e^{-c{\ell}^{d}}$ for some $c>0$. When the above event takes place, there can be at most $2\rho(2{\ell})^{d}/\left(C{\ell}-1\right)$ vertex-disjoint LR crossings of $\Lambda_{\ell}$ with length at least $C{\ell}$, for $C>1$. By taking $C$ large to assure $32^{d+1}\rho/C<c_{1}/$ 3 with $c_{1}$ as in Theorem 2.4, at cost to update $c_{1}$ and $c_{2}$ in (4), we get that Theorem 2.4 remains valid if we define $\mathcal{N}_{\ell}$ as the maximal number of vertex-disjoint LR crossings of $\Lambda_{\ell}$ with length upper bounded by $C{\ell}$ and included in the unique infinite cluster of $G_{d}$. We stress that, by Assumption 2.3(ii), any LR crossing has length at least 2/$2{\ell}/{\ell}_{*}$.
@@ 2.2 Transport via resistor networks
Having the resistor network $\text{RN}_{\ell}$, we consider the electrical potential $v_{\ell}:\mathcal{V}\cap S_{\ell}\to{\mathbb{R}}$ with value $1$ on $S_{\ell}^{-}$ and 00 on $S_{\ell}^{+}$. The potential $v_{\ell}$ is related to the current $\iota(x,y)$ that flows along the electric filament from $x$ to $y$ by the identity $\iota(x,y)=c_{x,y}(v_{\ell}(x)-v_{\ell}(y))$.
@@
We can finally present our second main result, concerning transport in the infinite cluster of our RCM. More precisely, we suppose that the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. Then, given $\rho\geq\lambda$, the graph $G_{d}(\omega,\underline{u})$ has a unique infinite cluster $\mathcal{C}(\omega,\underline{u})$ for $P_{\rho,\nu}\otimes Q$–a.a. $(\omega,\underline{u})$. We are interested in transport in the weighted undirected graph $\left(\mathcal{C}(\omega,\underline{u}),\mathcal{E}(\omega,\underline{u}),1\right)$, where the set $\mathcal{E}(\omega,\underline{u})$ is given by the edges in $G_{d}(\omega,\underline{u})$ between two points in the infinite cluster $\mathcal{C}(\omega,\underline{u})$.
@@
We recall that $D(\rho)$ describes the scaling limit of the direction conductivity of the infinite cluster with unit conductances. Indeed we can apply [Faggionato25]*Corollary 2.7 to this specific case (see Appendix A for some comments) getting the following.
@@
[…]
@@
By Theorem 2.4 and the Borel-Cantelli Lemma, for $P_{\rho,\nu}\otimes Q$–a.a. $(\omega,\underline{u})$, for $\ell$ large enough the cluster $\mathcal{C}(\omega,\underline{u})$ has at least $c_{1}{\ell}^{d-1}$ vertex-disjoint LR crossings of $\Lambda_{\ell}$. By Lemma 2.8 this implies, for $P_{\rho,\nu}\otimes Q$–a.a. $(\omega,\underline{u})$, that $\sigma_{\ell}(\omega,\underline{u})\geq c_{1}\frac{{\ell}^{2d-2}}{2\sharp(\hat{\omega}\cap\Lambda_{\ell})}$ for $\ell$ large. To control the denominator we observe that the ergodicity of the homogeneous PPP on ${\mathbb{R}}^{d}$ implies that $\lim_{\ell\to+\infty}\frac{\sharp(\hat{\omega}\cap\Lambda_{\ell})}{(2{\ell})^{d}}=\rho$ $P_{\rho,\nu}\otimes Q$–a.s. Hence, we conclude that $\liminf_{\ell\to+\infty}{\ell}^{2-d}\sigma_{\ell}(\omega,\underline{u})>0$ $P_{\rho,\nu}\otimes Q$–a.s. By combining this with Proposition 2.9 we get that $\kappa(\rho)>0$ and therefore $D(\rho)$ is strictly positive definite. ∎

### v2

@@ Abstract
We first study crossing statistics in random connection models (RCM) built on marked Poisson point processes on $\mathbb{R}^{d}$. Under general assumptions, we show exponential tail bounds for the number of crossings of a box contained in the infinite cluster for supercritical intensity of the point process, and percolation in slabs, in analogy with the Grimmett–-Marstrand theorem. We then present several applications to transport and diffusion phenomena. In particular, we prove the non-degeneracy of the effective homogenized matrix arising in the large-scale limit of random walks, exclusion processes, and resistor networks on the RCM, and the non-degeneracy of the effective diffusion constant for one-dimensional diffusion operators on the Euclidean graph associated with the RCM. As examples, we apply our results to Poisson–Boolean models and Mott variable range hopping random resistor network, providing a fundamental ingredient used in the derivation of Mott’s law.
@@ 1.1 Crossings and transport
Let us say that a random graph has a good crossing statistics if with high probability, for some $\kappa>0$, the subgraph given by the edges with conductances larger than $\kappa$ has at least order $\ell^{d-1}$ vertex-disjoint crossings of a box of side length $\ell$ in any given direction. Then, roughly, good crossing statistics ensure the non-degeneracy of $D$ in the 2-scale convergence approach of [Zhikov06, Faggionato23a] (at the basis of [Faggionato25, Faggionato22, FC]). For the supercritical Bernoulli percolation on ${\mathbb{Z}}^{d}$ (with unit conductances) the good crossing statistics were first proved by Kesten [Kesten82]*Chapter 11 and Chayes–Chayes [Chayes86]. Kesten’s proof is largely based on [Grimmett84] (cf. [Kesten82]*Remark (ii), p. 339). As presented in Section 6.3 the proof in [Chayes86] can be easily extended to get the same result for the supercritical Bernoulli percolation cluster itself. In particular, being a connected graph, the supercritical Bernoulli percolation cluster with unit or, more generally, positively lower bounded conductances is an effective graph to model transport in disordered conductors.
@@ 1.2 Main results
In this work we first aim at proving the above crossing property for more general random infinite clusters. We consider a random connection model (RCM) on a marked homogeneous Poisson point process (PPP) with density $\rho$ on ${\mathbb{R}}^{d}$, with i.i.d. marks, whose connection function $\varphi$, specifying the probability to have an edge between sites $x\not=y$, depends on yxy-x and on the marks of $x$ and $y$ (see Section 2 for a formal definition). When the connection function is an indicator (as in the Poisson–Boolean model), the insertion of the edge is deterministic and fully determined by the marked PPP. We point out that the RCM on a marked PPP is also widely used in the context of scale-free graphs (see e.g. [Gracar22]).
Under suitable assumptions on $\varphi$ (cf. Assumptions 2.3), we prove the following. Assume that, for some $\lambda>0$, the RCM has a unique infinite cluster for densities $\rho>\lambda$. Then (cf. Theorem 2.4), for all $\rho>\lambda$ there exist positive constants $c,c^{\prime}$ such that for all $\ell\geq 1$ the unique infinite cluster of the RCM has at least $c\ell^{d-1}$ left-right (LR) vertex-disjoint crossings of the box $\Lambda_{\ell}:=[-\ell,\ell]^{d}$ with probability at least $1-e^{-c^{\prime}\ell^{d-1}}$. Moreover, (cf. Theorem  2.6), under the same conditions, a sufficiently thick two-dimensional slab contains an infinite cluster, as in the classical result of Grimmett–Marstrand for Bernoulli percolation [Grimmett90]. In many applications, it is known that $\lambda$ can be taken equal to the critical density of the model. Moreover, in [Chebunin24], the uniqueness of the infinite cluster is reduced to checking a simple irreducibility condition. We also point out that Theorem 2.4 extends to a much larger class the results of [Faggionato21] obtained for RCM on marked PPP whose connection function is an indicator function satisfying somewhat restrictive monotonicity properties.
By Menger’s theorem, the maximal number of vertex-disjoint LR crossings corresponds to the minimal cardinality of vertex cutsets separating the two faces of the box $\Lambda_{\ell}$ under consideration (strictly speaking, the two half-strips, see Definition 2.2 below). The supercritical percolation cluster can be thought of as a capacity network where each edge has capacity one and Theorem 2.4 could be thought as a first step in investigating capacity networks built on marked RCMs. For both classical and more recent results about capacity networks we refer e.g. to [Grimmett84, Chayes86, Rossignol18, Zhang18] and references therein.
@@
1.3 Proof outline and comments on the strategy
We conclude with a rough outline of the proof strategy of the crossing result given by Theorem 2.4, emphasising some technical novelties and difficulties. We point out that this proof is the most demanding one and Theorem 2.6 emerges as a simple corollary of an intermediate result in this proof, while Theorems 2.112.133.13.73.8 follow relatively simply from Theorem 2.4 and known results, so we direct the reader to the respective full proofs.
Firstly, by a classical argument going back to [Chayes86, Grimmett84], it is enough to prove good crossing statistics in a quasi-two-dimensional setting. Then, LR crossings are constructed by studying a suitable growth process (see Section 6). It is related to the one introduced in [Tanemura93]*Section 4 for the blob model (i.e. the Poisson–Boolean model with a deterministic radius). However, we were led to extend and adapt Tanemura’s process, in order to recover crossings contained in the infinite cluster.
@@ 1.3 Proof outline and comments on the strategy
The above main difficulty of the models under consideration represented by ‘bad’ marks also impacts the dynamical renormalisation performed in Section 7 and illustrated in Figure 1. Indeed, the necessity to find ‘good’ marks progressively offsets renormalisation, imposing correcting the positioning at each step. Furthermore, preserving sufficient independence throughout the dynamical exploration in the continuum setting also requires some care in devising the renormalization.
Finally, let us comment on an alternative route in light of recent progress. Quantitative proofs of the Grimmett–Marstrand slab result have been obtained for Bernoulli percolation on ${\mathbb{Z}}^{d}$ [Duminil-Copin21] and on general transitive graphs with polynomial growth [Contreras24] (also see [Easo23] for faster growth), using dynamical and static renormalization, respectively. Our approach follows [Duminil-Copin21], using dynamical renormalization and an exploration algorithm adapted from Tanemura [Tanemura93] to establish stochastic domination between the explored region and the corresponding region in the supercritical Bernoulli percolation. Alternatively, one could have tried to extend the methods of [Contreras24] to RCMs. This approach is technically more involved but would yield additional information. In this case, the renormalized process would have finite-range dependence and, by the Liggett–Schonmann–Stacey theorem [Liggett97], would stochastically dominate a Bernoulli percolation that can be taken to be supercritical. By controlling local uniqueness events (see Definition 5.2) at the renormalisation scale, disjoint crossings in the renomalized process would lead to disjoint crossings in the original model. In particular, under this alternative approach, Tanemura’s construction would not be required. However, the difficulties arising from points with bad marks would remain.
@@ 2.1 Random connection model crossings
We let $J:=\{(i,j)\in{\mathbb{N}}_{+}\times{\mathbb{N}}_{+}\,:\,1\leq i<j\}$, where ${\mathbb{N}}_{+}:=\{1,2,\dots\}$, and denote by $\underline{u}=(u_{i,j})_{(i,j)\in J}$ a generic element on $0,1]^{J}$. We endow $0,1]^{J}$ with the product Lebesgue probability measure $Q=\bigotimes_{(i,j)\in J}\mathrm{d}x$.
Definition 2.1 (Graph $G_{d}$).
@@
[…]
@@
${\mathbb{S}}_{\ell}^{-}:=\{x\in{\mathbb{S}}_{\ell}\,:\,x_{1}<-{\ell}\}\,,\hskip 18.49988pt{\mathbb{S}}_{\ell}^{+}:=\{x\in{\mathbb{S}}_{\ell}\,:\,x_{1}>{\ell}\}\,,\hskip 18.49988pt\Lambda_{\ell}:=[-{\ell},{\ell}]^{d}\,.$
@@
Assumption 2.3(iv) is known to hold in several models. Chebunin and Last [Chebunin24] (see there for a detailed discussion on previous results) provide a very general criterion implying Assumption 2.3(iv). To state it we introduce, for $n\in{\mathbb{N}}_{+}$, the nn–th convolution of $\varphi$ given by the function $\varphi^{(n)}:({\mathbb{R}}^{d}\times{\mathbb{M}})^{2}\to[0,1]$ defined recursively by
$\varphi^{(1)}(t_{1},t_{2}):=\varphi(t_{1},t_{2})\hskip 18.49988pt\text{ and }\hskip 18.49988pt\varphi^{(n+1)}(t_{1},t_{2}):=\int\varphi^{(n)}(t_{1},t)\varphi(t,t_{2})(\lambda\mathcal{L}\otimes\nu)(dt)\,.$ (1)
@@
[…]
@@
Suppose the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. Then for any $\rho>\lambda$ there exist $c_{1},c_{2}>0$ such that, for all $\ell$ large enough, we have
@@
We call length of a LR crossing the number of its edges. By a large deviation estimate for Poisson random variables, for $\ell\geq 1$ the event $\sharp(\hat{\omega}\cap\Lambda_{\ell})\leq 2\rho(2\ell)^{d}$ has $P_{\rho,\nu}$–probability at least $1-e^{-c{\ell}^{d}}$ for some $c>0$. When the above event takes place, there can be at most $2\rho(2{\ell})^{d}/\left(C{\ell}-1\right)$ vertex-disjoint LR crossings of $\Lambda_{\ell}$ with length at least $C{\ell}$, for $C>1$. By taking $C$ large to assure $2^{d+1}\rho/C<c_{1}/3$ with $c_{1}$ as in Theorem 2.4, at cost to update $c_{1}$ and $c_{2}$ in (4), we get that Theorem 2.4 remains valid if we define $\mathcal{N}_{\ell}$ as the maximal number of vertex-disjoint LR crossings of $\Lambda_{\ell}$ with length upper bounded by $C{\ell}$ and included in the unique infinite cluster of $G_{d}$. We stress that, by Assumption 2.3(ii), any LR crossing has length at least $2{\ell}/{\ell}_{*}$.
The proof of Theorem 2.4 has the following notable byproduct (which is an immediate consequence of the Borel-Cantelli lemma and Eq. (37) in Section 7).
Theorem 2.6 (Percolation in slabs).
Suppose the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. Then for any $\rho>\lambda$ there exists $L>0$ such that the slab $\Sigma_{L}={\mathbb{R}}^{2}\times[-L,L]^{d-2}$ satisfies
$P_{\rho,\nu}\otimes Q(\Sigma_{L}\text{ contains an unbounded path of $G_{d}$})=1.$
Stating Theorem 2.6 explicitly in the present version was prompted by the posterior but independent work of Penrose [Penrose25], where Theorem 2.6 is proved in the particular case $\varphi:((x,m),(x^{\prime},m^{\prime}))\mapsto\phi(|x-x^{\prime}|_{p})$ for some $p\in[1,\infty]$ and nonincreasing $\phi:[0,\infty)\to[0,1]$ with finite support. [^1] The version in [Penrose25] is slightly different, stating that, for a thick enough slab, with positive probability, the origin (added to the PPP with density $\rho$ and marked independently with law $\nu$) has an infinite cluster in the RCM. By using Campbell’s identity (cf. [Daley88]), one can easily show the equivalence of the two formulations. The proof of [Penrose25] employs the original Grimmett–Marstrand approach, contrary to the present work. Previously, Last, Penrose and Zuyev [Last17a] proved Theorem 2.6 for the Poisson–Boolean model with bounded random radii, that is ${\mathbb{M}}=[0,1]$ and $\varphi((x,m),(x^{\prime},m^{\prime}))=\mathbbm{1}_{|x-x^{\prime}|_{2}\leq m+m^{\prime}}$. The result was also proved by Dembin and Tassion [DembinToappear] for unbounded radii under minimal moment assumptions (cf. Section 3.1).
@@ 2.2 Transport via resistor networks
Having the resistor network $\text{RN}_{\ell}$, we consider the electrical potential $v_{\ell}:\mathcal{V}\cap S_{\ell}\to{\mathbb{R}}$ with value $1$ on $S_{\ell}^{-}$ and 0 on $S_{\ell}^{+}$. The potential $v_{\ell}$ is related to the current $\iota(x,y)$ that flows along the electric filament from $x$ to $y$ by the identity $\iota(x,y)=c_{x,y}(v_{\ell}(x)-v_{\ell}(y))$.
@@
We can finally present our second main result, concerning transport in the infinite cluster of our RCM. More precisely, we suppose that the triple $(\lambda,\nu,\varphi)$ satisfies Assumptions 2.3. Then, given $\rho\geq\lambda$, the graph $G_{d}(\omega,\underline{u})$ has a unique infinite cluster $\mathcal{C}(\omega,\underline{u})$ for $P_{\rho,\nu}\otimes Q$–$a.$ $(\omega,\underline{u})$. We are interested in transport in the weighted undirected graph $\left(\mathcal{C}(\omega,\underline{u}),\mathcal{E}(\omega,\underline{u}),1\right)$, where the set $\mathcal{E}(\omega,\underline{u})$ is given by the edges in $G_{d}(\omega,\underline{u})$ between two points in the infinite cluster $\mathcal{C}(\omega,\underline{u})$.
@@
We recall that $D(\rho)$ describes the scaling limit of the direction conductivity of the infinite cluster with unit conductances. Indeed we can apply [Faggionato25]*Corollary 2.7 to this specific case (see Appendix A for some comments) to obtain the following.
@@
$\lim_{\ell\to+\infty}{(2\ell)}^{2-d}\sigma_{\ell}(\omega,\underline{u})=\rho\,\kappa(\rho)\hskip 18.49988ptP_{\rho,\nu}\otimes Q\text{--a.s.}$ (6)
@@
By Theorem 2.4 and the Borel-Cantelli Lemma, for $P_{\rho,\nu}\otimes Q$–$a.$ $(\omega,\underline{u})$, for $\ell$ large enough the cluster $\mathcal{C}(\omega,\underline{u})$ has at least $c_{1}{\ell}^{d-1}$ vertex-disjoint LR crossings of $\Lambda_{\ell}$. By Lemma 2.9 this implies, for $P_{\rho,\nu}\otimes Q$–$a.$ $(\omega,\underline{u})$, that $\sigma_{\ell}(\omega,\underline{u})\geq c_{1}\frac{{\ell}^{2d-2}}{2\sharp(\hat{\omega}\cap\Lambda_{\ell})}$ for $\ell$ large. To control the denominator we observe that the ergodicity of the homogeneous PPP on ${\mathbb{R}}^{d}$ implies that $\lim_{\ell\to+\infty}\frac{\sharp(\hat{\omega}\cap\Lambda_{\ell})}{(2{\ell})^{d}}=\rho$ $P_{\rho,\nu}\otimes Q$–a.s. Hence, we conclude that $\liminf_{\ell\to+\infty}{\ell}^{2-d}\sigma_{\ell}(\omega,\underline{u})>0$ $P_{\rho,\nu}\otimes Q$–a.s. By combining this with Proposition 2.10 we get that $\kappa(\rho)>0$ and therefore $D(\rho)$ is strictly positive definite. ∎